OpenStudy (thaliafortunx):

How many multiples of 5 are there between 199 and 1,198? Hint: an = a1 + d(n − 1), where a1 is the first term and d is the common difference.

OpenStudy (welshfella):

yes this is an arithmetic series where common difference is 5 The first term here would be 200 as its the next number after 199 divisible by 5

OpenStudy (thaliafortunx):

I get stuck at an=205+5n, I don't know if I'm doing the right steps

OpenStudy (welshfella):

sorry i got cut off. The equation is 1195 = 200 + (n - 1)d find n This will give you the answer

OpenStudy (jango_in_dtown):

1195=200+(n-1)*5 therefore 995=(n-1)*5 i.e. n-1=995/5 =199 i.e. n=199+1=200

OpenStudy (sshayer):